OpenStudy (anonymous):

Find the unit vector for each of the following in the direction of the given vector; verify that result has a magnitude of 1 a) u=<0,-2> b)w=-6i

ganeshie8 (ganeshie8):

find the magnitude of given vector, and divide

ganeshie8 (ganeshie8):

a) u=<0,-2> can u find the magnitude of this vector ?

OpenStudy (anonymous):

Don't know how to

ganeshie8 (ganeshie8):

magnitude of \(<x, y>\) is \(\sqrt{x^2 + y^2}\)

OpenStudy (anonymous):

so 2?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

yes, magnitude of <0, -2> is simply 2

ganeshie8 (ganeshie8):

so, the unit vector is \(\large <\frac{0}{2}, ~\frac{-2}{2}> = <0, ~-1> \)

OpenStudy (anonymous):

ohh.. That was so simple. Now what about b?

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

b)w=-6i

ganeshie8 (ganeshie8):

j component is missing, so write it as below : b)w=-6i + 0j

ganeshie8 (ganeshie8):

find its magnitude and divide

OpenStudy (anonymous):

okay, thank you. You made this beyond simple.

ganeshie8 (ganeshie8):

good to hear :) so wat did u get for b ?

OpenStudy (anonymous):

wait, would I find the magnitude of that the same way I did for a?

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

(6,0)?

ganeshie8 (ganeshie8):

b)w=-6i + 0j in vector form it is : <-6, 0>

OpenStudy (anonymous):

oh -6. I came close

ganeshie8 (ganeshie8):

:) find the mag

OpenStudy (anonymous):

6?

ganeshie8 (ganeshie8):

correct ! 6 is the magnitude to get the unit vector, divide the mag

ganeshie8 (ganeshie8):

unit vector = \(\large \frac{-6i}{6} = -i\)

OpenStudy (anonymous):

Thank you! :)

ganeshie8 (ganeshie8):

np :)